OpenStudy (anonymous):

1. Give the domain and range of this function:y=1(x+20)^2 2.Give the asymptotes and intercepts for this function: y=1(x-2)+2 3. identify the shifts: x=1/(x+4) -4 4. List all intervals (increasing, decreasing, positive, negative):y=(x^2)(x^2-2)-1/2 Any help is appreciated !!! thxs in Advance

OpenStudy (anonymous):

\(\large\color{midnightblue}{ y=(x+20)^2 }\) there is no restriction for the domain. i.e. you can plug any number for x, and this will not make a function undefined: Calculus wise, at any point, (for any x=a) 1) \(\large\color{midnightblue}{ f(a) }\) exisits 2) \(\large\color{midnightblue}{ \displaystyle\lim_{x \rightarrow a}f(x) }\) exist (saying that both sides of the limit are equivalent) 3) \(\large\color{midnightblue}{ \displaystyle\lim_{x \rightarrow a}f(x)=f(a) }\) domain: \(\large\color{midnightblue}{ (-\infty,\infty)}\) the range, gives us this: \(\large\color{midnightblue}{(x+20)^2}\) is always positive. so the least it can get, is when x=0. when x=-20. when x=-20, then (x+20)^2 is zero, but it won't be a negative, so: range: \(\large\color{midnightblue}{ [0,-\infty) }\)

OpenStudy (anonymous):

assuming a typo, it should be: \(\large\color{midnightblue}{ y=(x-2)^2+2 }\) for x it is a continous function, and y will be at least a 2.

OpenStudy (anonymous):

when x=0, the y=6 when y=0 the x=2\(\color{midniblue}{ \pm}\)\(\color{midniblue}{ i\sqrt{2}}\)

OpenStudy (anonymous):

(no real number y-intercept)

OpenStudy (anonymous):

next time I would ask one question at a time, and make no typos, you forgot the slashes... good luck

OpenStudy (anonymous):

thanks so much, how do you give a medal?

OpenStudy (anonymous):

so what to do with helping you

OpenStudy (anonymous):

I really need more help on number 2 Give the asmyptotes and intercepts of the function: y=1/(x-2)+2